OpenStudy (anonymous):

Find sin Θ if tan Θ= the square root of 2 and falls in Quadrant 3. HELP PLEASE ASAP @robtobey @radar

OpenStudy (anonymous):

tan (theta) = sqrt(2) = sqrt(2) / 1 Recall that tan (theta) = y/x and sin (theta) = y/r In this case, opposite side = sqrt (2) and adjacent side is 1. The fact that theta is in quadrant 3 also implys that both x and y values are negative Thus x = -1 and y = -sqrt(2) Now simply use Pythagorean theorem to calculate r and write out x/r for sin(theta)

OpenStudy (anonymous):

so the answer is -1/-sqrt(2) ? @CLOUDYskies

OpenStudy (mertsj):



OpenStudy (mertsj):

Use the Pythagorean Theorem to find the hypotenuse.

OpenStudy (mertsj):